OpenStudy (anonymous):

Will medal!!! 3(x + 1) = -2(x - 1) + 6

OpenStudy (anonymous):

are e solving for x?

OpenStudy (anonymous):

we*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply the distributive property of multiplication over addition, so we can write this: \[3x + 3 = - 2x + 2 + 6\]

OpenStudy (anonymous):

is it 1?

OpenStudy (michele_laino):

please keep in mind that: \(-2(x-1)=-2x-2(-1)=-2x+2\)

OpenStudy (michele_laino):

correct! the solution is \(x=1\)

OpenStudy (anonymous):

Thank you!!!!!

OpenStudy (michele_laino):

:)